Show that ¬ (P"\iff"Q)"\iff"(P V Q) Λ ¬(P Λ Q) "\iff"(P Λ ¬Q) V (¬ P Λ Q) without using truth table


1
Expert's answer
2022-01-31T12:32:02-0500

Since

"\neg \left( {P \leftrightarrow Q} \right) \equiv \neg \left( {\left( {P \to Q} \right) \wedge \left( {Q \to P} \right)} \right) \equiv \neg \left( {\left( {\neg P \vee Q} \right) \wedge \left( {\neg Q \vee P} \right)} \right) \equiv \neg \left( {\neg P \vee Q} \right) \vee \neg \left( {\neg Q \vee P} \right) \equiv \left( {P \wedge \neg Q} \right) \vee \left( {Q \wedge \neg P} \right)"

Then

"\neg \left( {P \leftrightarrow Q} \right) \leftrightarrow \left( {P \wedge \neg Q} \right) \vee \left( {\neg P \wedge Q} \right)"

Since

"\left( {P \wedge \neg Q} \right) \vee \left( {\neg P \wedge Q} \right) \equiv \left( {P \vee \neg P} \right) \wedge \left( {\neg Q \vee \neg P} \right) \wedge \left( {P \vee Q} \right) \wedge \left( {\neg Q \vee Q} \right) \equiv 1 \wedge \left( {\neg Q \vee \neg P} \right) \wedge \left( {P \vee Q} \right) \wedge 1 \equiv \left( {\neg Q \vee \neg P} \right) \wedge \left( {P \vee Q} \right) \equiv \left( {P \vee Q} \right) \wedge \neg \left( {P \wedge Q} \right)"

Then

"\left( {P \wedge \neg Q} \right) \vee \left( {\neg P \wedge Q} \right) \leftrightarrow \left( {P \vee Q} \right) \wedge \neg \left( {P \wedge Q} \right)"

But then

"\neg \left( {P \leftrightarrow Q} \right) \leftrightarrow \left( {P \vee Q} \right) \wedge \neg \left( {P \wedge Q} \right) \leftrightarrow \left( {P \wedge \neg Q} \right) \vee \left( {\neg P \wedge Q} \right)"

Q. E. D.